A boeing KC-135A airplane has a wingspan of 39.9 m and flies at a constant altitude in a northerly direction with a speed of 750 km/h. If the vertical component of the Earth's magnetic field is 5.2*10^-6 T, and its horizontal component is 1.5*10^-6 T, what is the induced emf between the wing tips?
Answer #1

$$ \begin{aligned} \text { Net Emf } &=V \times l \times B y \\ &=\frac{750}{3.6} \times 39.9 \times 5.2 \times 10^{-6} \\ &=4.32 \times 10^{-2} \mathrm{~V} . \end{aligned} $$